# Accelerating the Search for Superconductors Using Machine Learning

### v1

@@ III.2 Machine Learning Workflow
Specifically, we first classify whether a material is a superconductor or a non-superconductor. If a compound is predicted to be a superconductor, we apply regression-based machine learning models to estimate its critical temperature. Otherwise, we assign it a critical temperature of 0 K. This two-step approach enhances the reliability of our machine learning predictions. Fig. 2 provides a graphical representation of our machine learning workflow.
@@ IV Results
The SuperCon-MTG database hosts 13,731 compounds, including both reported superconductors and non superconducting compounds, after the data cleaning with algorithm described in Sec III.1. The cleaned database contains 1,584 non-superconducting compounds with a critical temperature of 0 K. The highest reported critical temperature in the dataset is 143 K for $\text{Hg}_{0.66}\text{Pb}_{0.34}\text{Ba}_{2}\text{Ca}_{1.98}\text{Cu}_{2.97}\text{O}_{8.4}$ [48]. The average critical temperature of the compounds in the dataset is 21.2 K. The first quartile ($Q_{1}$) lies below 2.05 K, meaning that 25% of compounds in the database have critical temperature $<$ 2.05 K. The second quartile ($Q_{2}$) is below 7.5 K and third quartile ($Q_{3}$) falls below 29.6 K.
@@
A clear class imbalance exists between superconductors and non-superconductors, as the number of superconducting materials is significantly higher. This imbalance arises due to limited availability of labeled non-superconducting materials data in materials science. Despite the existence of well-maintained theoretical and experimental databases such as Materials Project [34] which hosts over 150,000 inorganic materials, The Open Quantum Materials Database (OQMD) [49], which contains over 1.2 million DFT calculated thermodynamic and structural properties and the Inorganic Crystal Structure Database (ICSD) [50] hosting more than 300,000 experimetally determined compounds, none of these databases explicitly label materials as superconducting / non-superconducting. Due to this limitation, we cannot manually add materials from these datasets to balance the classes, which makes it difficult to ensure a well-distributed dataset for machine learning models.
@@
Figure 3: (a) Histogram of $T_{c}$ in SuperCon-MTG, where the first bin (at $T=0$ K) has a width of 1 K, while all other bins have a width of 3 K with inset of distribution of $T_{c}$ for 5 largest classes of superconductors. (b) The distribution of compounds in the SuperCon-MTG database by class, including Cuprates, Alloys, Iron Chalcogenides, Iron Pnictides, Heavy-Fermion materials, Silicides, Borides, Oxides, Borocarbides, Tellurides, Germanides, Nitrides, Transition Metal Pnictides, Intercalated Graphite, Hydrides, Bismuthates, Elemental Superconductors (labeled as C2), A15 Compounds, Fullerides (labeled as C1), $\text{BiS}_{2}$-based materials, other Iron-based Compounds, Chevrel phases (labeled as C3), Dichalcogenides (labeled as C4), and Unclassified materials.
@@
We classify the compounds in SuperCon-MTG into twenty-four classes based on their chemical composition. Cuprates (copper oxide compounds) constitute the largest fraction, accounting for 34.3% of the database, followed by alloys at 20%. Iron chalcogenides make up approximately 5.6%, Iron pnictides 4.8%, and Heavy Fermion compounds around 4.4%. Unidentified class of compounds fall under ‘Unclassified’ constituting 4.6%. The remaining 18 classes collectively contribute 26.3% to the database. The extensive research on cuprates, driven by their potential as high $T_{c}$ superconductors, is evident from their dominant presence, making up nearly one-third of the dataset.
@@ IV.1 Unsupervised Learning: Principal Component Analysis
Previous studies have demonstrated the clustering of compounds in the $T_{c}$ vs. $\overline{N_{v}}$ space [39, 40]. This clustering arises because certain classes of materials exhibit superconducting critical temperatures within specific ranges and tend to share common crystal structures. As a result, clustering in this space is a natural consequence rather than an insightful distinction. In our analysis, we extend this idea by projecting superconducting materials into two principal components using Principal Component Analysis (PCA). We observe distinct clustering patterns, with compounds grouping according to their structural and chemical similarities. Notably, cuprates, elemental superconductors, and carbon-based compounds exhibit clear clustering. Other classes also cluster, though with some degree of overlap between different groups, making their boundary diffused. Clustering of these compounds in PCA space can be seen as in Fig. 4.
@@ IV.2 Hyperparameter Tuning
Random Forest [51] is an ensemble learning method that constructs multiple decision trees using bagging. Specifically, it selects random samples with replacement (bootstrapping) from the dataset to train each tree. This means that after a sample is drawn from the dataset, it is returned to the bag, making it possible for the same sample to be selected multiple times for training different trees. For regression tasks, the final output is the average prediction of individual trees, whereas for classification, it is determined by majority voting. This random sampling process helps reduce overfitting and variance.
The Random Forest model consistently outperforms both XGBoost [52] and Extra Trees for classification and regression tasks related to superconductivity prediction (see Supplementary Information Sec. II [41] for details). The overall workflow is outlined in Sec III.2.
@@ IV.2.1 Classification
We perform hyperparameter tuning to select the best parameters for fitting our Random Forest classification model using Bayesian optimization with the help of the scikit-optimize [53] library in Python. Unlike random or exhaustive search methods such as Random Search and Grid Search, Bayesian optimization is based on a probabilistic model to efficiently explore the hyperparameter space. We perform 400 iterations to search for the optimal hyperparameters. The selected hyperparameters are detailed in Supplementary Information (SI), Sec. III A [41].
@@ IV.2.2 Regression
We select the predicted superconductors using the Random Forest classification model, as described in Section IV.2.1. The training and test datasets now consist of 9,800 and 2,584 predicted superconductors, respectively. Bayesian optimization is performed to obtain the best hyperparameters for the regression model used to predict the critical temperature using the Random Forest algorithm. The selected hyperparameters are detailed in Supplementary Information (SI), Sec. III B [41] .
@@
It is clear (see Fig. 6) that the ML model learns effectively and performs well on the training dataset. For the test dataset, the model makes accurate predictions but tends to underfit for $T_{c}$ $>$ 40 K. There are abrupt predictions at 0 K, which primarily correspond to non-superconductors misclassified as superconductors by the Random Forest classification model. Most of these compounds have compositions near the boundary between superconducting and non-superconducting phases or they belong to the non-superconductors imputed into the database reported in [47]. We expect these compounds to be superconducting upon tuning of chemical composition or physical parameters like pressure.
@@ IV.3 SHapley Additive exPlanations (SHAP) Analysis of the Random Forest Model
Random Forest is a black-box model, making it difficult to interpret the decision-making process behind its predictions. Explainable AI (XAI) is a field in artificial intelligence that aims to make models more interpretable and explainable. Some of the most commonly used XAI techniques are SHAP [33] and LIME [54].
SHAP explains machine learning model predictions using ideas from game theory. It fairly distributes the contribution of each feature to the final predictions by calculating Shapley values [55], which represent each feature’s average marginal contribution across all possible coalitions. A higher Shapley (SHAP) value indicates that the feature increases the predicted value by that amount relative to the model’s baseline value, which is the average of the predicted value over the training dataset. The baseline value for our model in 23.6 K.
@@
In Fig. 7, we present a bar plot of the mean absolute SHAP values for each feature, which indicates the ranking of feature importance. A higher mean absolute SHAP value signifies a greater impact on the model’s predictions but does not indicate whether an increase in the feature value leads to a positive or negative SHAP value. ‘Maximum orbital radius’ is the most important feature driving critical temperature prediction, followed by ‘Median electronegativity’, while the last bar represents the contribution of the remaining 21 features.
In Fig. 8, we present feature distributions displayed as violin plot illustrating the relationship between feature values and SHAP values. The feature ‘Maxima Orbital Radius’, which has the most significant influence on critical temperature, tends to decrease the critical temperature from the baseline value as the feature value increases. Conversely, ‘Median Orbital Radius’ exhibits a negative SHAP value for higher feature values and a positive SHAP value for lower feature values, meaning that lower values of this feature increase the predicted critical temperature above the baseline, while higher values decrease it below the baseline.
Figure 8: Violin plot of feature values and SHAP values.
@@ IV.4 Random Forest model performance with 5 Key descriptors
A Random Forest regression model was trained for predicting critical temperature with these five key features for materials classified as superconductors by Random Forest classification model in Section IV.2.1. Hyperparameters for Random forest regression model with 5 Key features are obtained by performing Bayesian Optimization and selected parameters are detailed in Supplementary Information (SI), Sec. III C [41].
@@
RMSE (K) 3.2 6.2
MAE (K) 5.9 11.3
@@ IV.5 Classwise performance of Machine Learning Model
Around 12,384 compounds are classified as superconductors in SuperCon-MTG using Random Forest classification model in Sec IV.2.1 . These materials fall into 24 distinct classes based on the chemical composition. We train a random forest regression model to predict critical temperature for each class of these compounds, to improve the reliability of ML model predictions, we merge smaller classes with fewer than 170 entries, ensuring adequate dataset size for training. More details on the statistics of classes of compounds can be found in Supplementary Information (SI), Section IV [41].
@@
The $R^{2}$ score and normalized error, defined as the ratio of the root mean squared error to the average critical temperature of each class, are as shown in Fig. 10. We observe that the random forest model predicts the critical temperature reasonably well for Bi-based materials [56] and borocarbides [57], both of which are potentially unconventional superconductors. Similarly, it performs well for transition metal pnictides [58] and borides [59], which are considered conventional superconductors. On the other hand, the accuracy of model predictions for $T_{c}$’s of other classes of superconductors such as cuprates, nitrides, oxides, and iron-based pnictides and chalcogenides, are relatively lower. This indicates that the class-wise trained model performs well for classes where compositional trends dominate, while its performance slightly declines for classes where crystal structure may play a more significant role. Overall, the model shows reasonably good performance and generalizes $T_{c}$ prediction effectively, when trained across classes using only composition-based features.
@@
Figure 10: Performance metrics for regression using a Random Forest model (with default parameters in scikit-learn [60]) across different classes of superconducting materials.
V Search for Superconducting Materials
V.1 Superconductors from the Literature
@@
In the previous section, we demonstrated that our machine learning model is reasonably robust in classifying materials as superconductors and predicting their $T_{c}$. To further validate its performance beyond the SuperCon-MTG dataset, we evaluate materials reported after 2016 that are not included in the training or test datasets. We select 27 compounds that are known to exhibit superconductivity. Our model successfully classifies all of these 27 materials as superconductors and provides reasonably accurate critical temperature predictions. The results are presented in Table 4. The $\text{Model}_{30}$ and $\text{Model}_{5}$ columns represent $T_{c}$ predictions using machine learning models with 30 and 5 descriptors, respectively.
@@
Compound $\text{Model}_{30}$(K) $\text{Model}_{5}$(K) Exp $T_{c}$(K) Class $\text{BaB}_{5}$ 16.3 17.6 16.3 [61] Borides RhSeTe 4.8 4.1 4.72 [62] Tellurides $\text{Re}_{7}\text{Nb}_{3}$ 5.4 6 5.2 [63] Alloys $\text{ZrTe}_{1.8}$ 2.8 3.7 3.2 [64] Tellurides $\text{Re}_{7}\text{Ti}_{3}$ 4.4 5.3 3.8 [63] Alloys TiIrGe 3 3.4 2.24 [65] Germanides $\text{Na}_{2}\text{CoSe}_{2}\text{O}$ 7 5.5 6.3 [66] Oxides $\text{Re}_{7}\text{Zr}_{3}$ 4.8 4.6 5.6 [63] Alloys $\text{Pt}_{1.8}\text{Ir}_{1.2}\text{Zr}_{5}$ 4.5 3 3.2 [67] Alloys $\text{Re}_{7}\text{Hf}_{3}$ 4.7 4 6 [63] Alloys $\text{Re}_{7}\text{Ta}_{3}$ 5 5.5 3.6 [63] Alloys $\text{TaIr}_{2}\text{B}_{2}$ 3.6 2.7 5.2 [68] Borides $\text{Pt}_{1.5}\text{Ir}_{1.5}\text{Zr}_{5}$ 4.9 4.9 3.2 [67] Alloys $\text{Pt}_{1.2}\text{Ir}_{1.8}\text{Zr}_{5}$ 4.5 5.1 2.8 [67] Alloys $\text{Cu}_{2}\text{Si}$ 2.1 2.7 4.1 [69] Silicides HfRhGe 3.8 4 1.7 [70] Germanides $\text{Pt}_{2.7}\text{Ir}_{0.3}\text{Zr}_{5}$ 6 2.5 3.8 [67] Alloys $\text{Mo}_{20}\text{Re}_{6}\text{Si}_{4}$ 7.7 12.3 5.4 [28] Silicides $\text{Pt}_{2.1}\text{Ir}_{0.9}\text{Zr}_{5}$ 5.8 3.5 3.4 [67] Alloys $\text{Pt}_{0.9}\text{Ir}_{2.1}\text{Zr}_{5}$ 5.1 5.5 2.7 [67] Alloys $\text{Pt}_{2.4}\text{Ir}_{0.6}\text{Zr}_{5}$ 6.2 3.9 3.6 [67] Alloys HfIrGe 3.1 3.5 5.64 [65] Germanides $\text{Pt}_{0.45}\text{Ir}_{2.55}\text{Zr}_{5}$ 5.3 5.9 2.6 [67] Alloys $\text{Pt}_{0.6}\text{Ir}_{2.4}\text{Zr}_{5}$ 5.3 5.9 2.6 [67] Alloys $\text{NbIr}_{2}\text{B}_{2}$ 4.2 3.2 7.2 [68] Borides $\text{Pt}_{3}\text{Zr}_{5}$ 3.5 2.8 7.3 [71] Alloys $\text{Re}_{8}\text{NbTa}$ 3.4 3.9 7.7[72] Alloys ReTc 6.9 6.3 11.5 [31] [^1] DFT predicted $T_{c}$ Alloys
V.2 Using the compounds in Materials Project

### v2

@@ III.2 Machine Learning Workflow
The SuperCon-MTG dataset used for training the model is split in an 80–20 ratio for training and testing, respectively (see Supplementary Information Sec. III [SI] for details). Specifically, we first classify whether a material is a superconductor or a non-superconductor. If a compound is predicted to be a superconductor, we apply regression-based machine learning models to estimate its critical temperature. Otherwise, we assign it a critical temperature of 0 K. This two-step approach enhances the reliability of our machine learning predictions. Fig. 2 provides a graphical representation of our machine learning workflow.
@@ IV Results
After data cleaning with the algorithm described in Sec. III.1, the SuperCon-MTG dataset hosts 13,731 compounds, including both reported superconductors and non-superconducting compounds. The cleaned database contains 1,584 non-superconducting compounds with a critical temperature of 0 K. The highest reported critical temperature in the dataset is 143 K for $\text{Hg}_{0.66}\text{Pb}_{0.34}\text{Ba}_{2}\text{Ca}_{1.98}\text{Cu}_{2.97}\text{O}_{8.4}$ [SHAO1995]. The average critical temperature of the compounds in the dataset is 21.2 K. The first quartile ($Q_{1}$) lies below 2.05 K, meaning that 25% of compounds in the database have critical temperature << 2.05 K. The second quartile ($Q_{2}$) is below 7.5 K and third quartile ($Q_{3}$) falls below 29.6 K.
@@
A clear class imbalance exists between superconductors and non-superconductors, as the number of superconducting materials is significantly higher. This imbalance arises due to limited availability of labeled non-superconducting materials data in materials science. Despite the existence of well-maintained theoretical and experimental databases such as Materials Project [Jain2013] which hosts over 150,000 inorganic materials, The Open Quantum Materials Database (OQMD) [Saal2013], which contains over 1.2 million DFT calculated thermodynamic and structural properties and the Inorganic Crystal Structure Database (ICSD) [Belsky2002] hosting more than 300,000 experimetally determined compounds, none of these databases explicitly label materials as superconducting / non-superconducting. Due to this limitation, we cannot manually add materials from these databases to balance the classes, which makes it difficult to ensure a well-distributed dataset for machine learning models.
@@
Figure 3: (a) Histogram of $T_{c}$ in SuperCon-MTG, where the first bin (at $T=0$ K) has a width of 1 K, while all other bins have a width of 3 K with inset of distribution of $T_{c}$ for 5 largest classes of superconductors. (b) The distribution of compounds in the SuperCon-MTG dataset by class, including Cuprates, Alloys, Iron Chalcogenides, Iron Pnictides, Heavy-Fermion materials, Silicides, Borides, Oxides, Borocarbides, Tellurides, Germanides, Nitrides, Transition Metal Pnictides, Intercalated Graphite, Hydrides, Bismuthates, Elemental Superconductors (labeled as C2), A15 Compounds, Fullerides (labeled as C1), $\text{BiS}_{2}$-based materials, other Iron-based Compounds, Chevrel phases (labeled as C3), Dichalcogenides (labeled as C4), and Unclassified materials.
@@
We classify the compounds in SuperCon-MTG into twenty-four classes based on their chemical composition. Cuprates (copper oxide compounds) constitute the largest fraction, accounting for 34.3% of the database, followed by alloys at 20%. Iron chalcogenides make up approximately 5.6%, Iron pnictides 4.8%, and Heavy Fermion compounds around 4.4%. Unidentified class of compounds fall under ‘Unclassified’ constituting 4.6%. The remaining 18 classes collectively contribute 26.3% to the dataset. The extensive research on cuprates, driven by their potential as high $T_{c}$ superconductors, is evident from their dominant presence, making up nearly one-third of the dataset.
@@ IV.1 Unsupervised Learning: Principal Component Analysis
Previous studies have demonstrated the clustering of compounds in the $T_{c}$ vs. $\overline{N_{v}}$ space [Villars1988, Rabe1992]. This clustering arises because certain classes of materials exhibit superconducting critical temperatures within specific ranges and tend to share common crystal structures. As a result, clustering in this space is a natural consequence rather than an insightful distinction. In our analysis, we extend this idea by projecting superconducting materials into two principal components using Principal Component Analysis (PCA). Two of the three QSD-based descriptors, weighted average of valence electron number and weighted average of orbital radius difference are the 9th and 10th most important features influencing the PC-1 component. We observe distinct clustering patterns, with compounds grouping according to their structural and chemical similarities. Notably, cuprates, elemental superconductors, and carbon-based compounds exhibit clear clustering. Other classes also cluster, though with some degree of overlap between different groups, making their boundary diffused. Clustering of these compounds in PCA space can be seen as in Fig. 4.
@@ IV.2 Hyperparameter Tuning
Random forest [Breiman2001] is an ensemble learning method that constructs multiple decision trees using bagging. Specifically, it selects random samples with replacement (bootstrapping) from the dataset to train each tree. This means that after a sample is drawn from the dataset, it is returned to the bag, making it possible for the same sample to be selected multiple times for training different trees. For regression tasks, the final output is the average prediction of individual trees, whereas for classification, it is determined by majority voting. This random sampling process helps reduce overfitting and variance.
The Random forest model consistently outperforms both XGBoost [Chen2016] and Extra Trees for classification and regression tasks related to superconductivity prediction (see Supplementary Information Sec. III [SI] for details). The overall workflow is outlined in Sec III.2.
@@ IV.2.1 Classification
We perform hyperparameter tuning to select the best parameters for fitting our Random forest classification model using Bayesian optimization with the help of the scikit-optimize [scikit-optimize] library in Python. Unlike random search or exhaustive methods such as grid search, Bayesian optimization is based on a probabilistic model to efficiently explore the hyperparameter space. We perform 400 iterations to search for the optimal hyperparameters. The selected hyperparameters are detailed in Supplementary Information (SI), Sec. IV A [SI].
@@ IV.2.2 Regression
We select the predicted superconductors using the Random forest classification model, as described in Section IV.2.1. The training and test datasets now consist of 9,800 and 2,584 predicted superconductors, respectively. Bayesian optimization is performed to obtain the best hyperparameters for the regression model used to predict the critical temperature using the Random forest algorithm. The selected hyperparameters are detailed in Supplementary Information (SI), Sec. IV B [SI].
@@
It is clear (see Fig. 6) that the ML model learns effectively and performs well on the training dataset. For the test dataset, the model makes accurate predictions but tends to underfit for $T_{c}$ >> 40 K. There are abrupt predictions at 0 K, which primarily correspond to non-superconductors misclassified as superconductors by the Random forest classification model. Most of these compounds have compositions near the boundary between superconducting and non-superconducting phases or they belong to the non-superconductors imputed into the dataset reported in [Hosono2015]. We expect these compounds to be superconducting upon tuning of chemical composition or physical parameters like pressure.
@@ IV.3 SHapley Additive exPlanations (SHAP) Analysis of the Random forest Model
Random forest is a black-box model, making it difficult to interpret the decision-making process behind its predictions. Explainable AI (XAI) is a field in artificial intelligence that aims to make models more interpretable and explainable. Some of the most commonly used XAI techniques are SHAP [SHAP] and LIME [LIME].
SHAP explains machine learning model predictions using ideas from game theory. It fairly distributes the contribution of each feature to the final predictions by calculating Shapley values [Shapley1951], which represent each feature’s average marginal contribution across all possible coalitions. A higher Shapley (SHAP) value indicates that the feature increases the model’s prediction relative to the baseline, which is the average of the predicted value over the training dataset. The baseline value for our model in 23.6 K.
@@
In Fig. 7, we present a bar plot of the mean absolute SHAP values for each feature, which indicates the ranking of feature importance. A higher mean absolute SHAP value signifies a greater impact on the model’s predictions. However, it does not indicate whether an increase in the feature value leads to a positive or negative SHAP value. ‘Maximum orbital radius’ is the most important feature driving critical temperature prediction, followed by ‘Median electronegativity’, while the last bar represents the contribution of the remaining 21 features.
In Fig. 8, we present feature distributions displayed as violin plot illustrating the relationship between feature values and SHAP values. The feature ‘Maxima Orbital Radius’, which has the most significant influence on critical temperature, tends to decrease the critical temperature from the baseline value as the feature value increases. Conversely, ‘Median Orbital Radius’ exhibits a negative SHAP value for higher feature values and a positive SHAP value for lower feature values. This indicates that lower values of this feature increase the predicted critical temperature above the baseline, while higher values decrease it below the baseline.
In the specific case of the ”maximum orbital radius” descriptor, this trend aligns with physical insights from copper-oxide based superconductors. A lower orbital radius means a smaller spatial extent of valence orbitals and weaker orbital overlap. This results in the reduced hopping amplitudes and lower bandwidth, and hence stronger electronic correlation. Similarly, for many conventional superconductors, larger orbital radii result in greater overlap and wider, more dispersed bands, reducing the electronic density of states at the Fermi level, which—via the Eliashberg relation—can lead to lower $T_{c}$. However, this trend could also reflect the fact that more dispersed bands lead to stronger electron–phonon coupling and higher-frequency phonons, both of which have a positive effect on $T_{c}$. Thus, the feature ”maximum orbital radius” has a mixed effect on the conventional class of superconductors. This indicates that the same feature behaves differently across various classes of superconductors. Consequently, the SHAP trend for this feature is difficult to interpret in a physically consistent way across the full dataset.
Additionally, average electronegativity has been shown to correlate with $T_{c}$ [Asokamani1989, Jayaprakash1993]. For example, copper-oxide-based superconductors typically exhibit electronegativity values between 2.12 and 2.53, and our SHAP analysis reveals its positive contributions to $T_{c}$ within this range. In contrast, elemental superconductors often lie in the 1.3–1.9 range, where SHAP values are predominantly negative, indicating suppression of $T_{c}$.
Thus, physically interpretable connections between the statistical descriptors and underlying superconducting mechanisms based on SHAP analysis reinforces the value of combining data-driven approaches to obtain insights for physical understanding.
@@ IV.4 Random forest model performance with 5 Key descriptors
A Random forest regression model was trained for predicting critical temperature with these five key features for materials classified as superconductors by Random forest classification model in Section IV.2.1. Hyperparameters for Random forest regression model with 5 Key features are obtained by performing Bayesian Optimization and selected parameters are detailed in Supplementary Information (SI), Sec. IV C [SI].
@@
MAE (K) 3.2 6.2
RMSE (K) 5.9 11.3
Figure 8: Violin plot of feature values and SHAP values.
@@ IV.5 Classwise performance of Machine Learning Model
Around 12,384 compounds are classified as superconductors in SuperCon-MTG using Random forest classification model in Sec IV.2.1 . These materials fall into 24 distinct classes based on the chemical composition. We train a Random forest regression model to predict critical temperature for each class of these compounds, to improve the reliability of ML model predictions, we merge smaller classes with fewer than 170 entries, ensuring adequate dataset size for training. More details on the statistics of classes of compounds can be found in Supplementary Information (SI), Section V [SI].
@@
The $R^{2}$ score and normalized error, defined as the ratio of the root mean squared error to the average critical temperature of each class, are as shown in Fig. 10. We observe that the Random forest model predicts the critical temperature reasonably well for Bi-based materials [Sleight2015] and borocarbides [Mazumdar2015], both of which are potentially unconventional superconductors. Similarly, it performs well for transition metal pnictides [Klemm2015] and borides [Chen2024], which are considered conventional superconductors. On the other hand, the accuracy of model predictions for $T_{c}$’s of other classes of superconductors such as cuprates, nitrides, oxides, and iron-based pnictides and chalcogenides, are relatively lower. In case of cuprate superconductors, the optimal Cu–O bond length is around 1.9 Å, and deviations from this value are known to suppress superconductivity. Similarly, the apical Cu–O bond length has been shown to correlate with $T_{c}$ [KAMBE2000]. In Iron-based materials, structural features such as the anion height (the vertical distance between the Fe atom and the anion plane) plays a critical role in superconductivity. An optimal anion height of approximately 1.38 Å has been empirically linked to the highest $T_{c}$ values in many FeAs-based superconductors [Chun2014]. For example, in the 1111 phase, replacing La with smaller rare-earth elements such as Nd or Sm increases the anion height from 1.33 Å to 1.38 Å and raises $T_{c}$ from 26 K to 56 K. However, increasing the anion height beyond this optimal value leads to a decrease in $T_{c}$, indicating a non-monotonic relationship. This effect is structurally linked to the Fe–pnictogen/chalcogen bond length, which influences both the anion height and the Fe–Pn/Ch–Fe bond angles. The highest $T_{c}$ values are observed when the FePn4 tetrahedron approaches ideal geometry, with bond angles close to $109.47^{\circ}$. This structural configuration optimizes orbital hybridization and supports strong spin fluctuation–mediated pairing, which is believed to be responsible for superconductivity in these materials. This indicates that the class-wise trained model performs well for classes where compositional trends dominate, while its performance slightly declines for classes where crystal structure may play a more significant role. Moreover, composition-only features do not capture carrier concentration, magnetic ordering, or phonon properties that do influence superconductivity. Previous machine-learning studies [Zhangr12020, Zhangr12022] have achieved reasonably realistic $T_{c}$ predictions by incorporating synthesis-specific parameters—such as sintering temperature, time, or pulsed-layer-deposition conditions. Overall, our model shows reasonably good performance and generalizes $T_{c}$ prediction effectively, when trained across classes using only composition-based features. Incorporating structure- and synthesis- specific features among descriptors as variables of predictive models is a promising direction for future work and may significantly improve the model performance, particularly in the high-$T_{c}$ predictions.
@@
Figure 10: Performance metrics for regression using a Random forest model (with default parameters in scikit-learn [scikit-learn]) across different classes of superconducting materials.
V Validation of our model and Search for new Superconductors
V.1 Superconductors from the recent literature
@@
In the previous section, we demonstrated that our machine learning model is reasonably robust in classifying materials as superconductors and predicting their $T_{c}$. To further validate its performance beyond the SuperCon-MTG dataset, we evaluate materials reported after 2016 that are not included in the training or test datasets. We select 27 compounds that are known to exhibit superconductivity. Our model successfully classifies all of these 27 materials as superconductors and provides reasonably accurate critical temperature predictions. The results are presented in Table 4. The $\text{Model}_{30}$ and $\text{Model}_{5}$ columns represent $T_{c}$ predictions using machine learning models with 30 and 5 descriptors, respectively. Similarly, we evaluate our model on 31 compounds previously reported as potential superconductors in earlier machine-learning works [Jung2024, Pereti2023, Gashmard2024, Kaplan2025]. This benchmarking demonstrates that our model attains performance on par with these prior approaches while employing a smaller, QSD-inspired descriptor set (see Supplementary Information Sec. VIII [SI] for details).
@@
Compound $\text{Model}_{30}$(K) $\text{Model}_{5}$(K) Exp $T_{c}$(K) Class $\text{BaB}_{5}$ 16.3 17.6 16.3 [Xie2023] Borides RhSeTe 4.8 4.1 4.72 [Patra2024] Tellurides $\text{Re}_{7}\text{Nb}_{3}$ 5.4 6 5.2 [Kushwaha2024] Alloys $\text{ZrTe}_{1.8}$ 2.8 3.7 3.2 [Correa2023] Tellurides $\text{Re}_{7}\text{Ti}_{3}$ 4.4 5.3 3.8 [Kushwaha2024] Alloys TiIrGe 3 3.4 2.24 [Meena2025] Germanides $\text{Na}_{2}\text{CoSe}_{2}\text{O}$ 7 5.5 6.3 [Cheng2024] Oxides $\text{Re}_{7}\text{Zr}_{3}$ 4.8 4.6 5.6 [Kushwaha2024] Alloys $\text{Pt}_{1.8}\text{Ir}_{1.2}\text{Zr}_{5}$ 4.5 3 3.2 [Watanabe2023] Alloys $\text{Re}_{7}\text{Hf}_{3}$ 4.7 4 6 [Kushwaha2024] Alloys $\text{Re}_{7}\text{Ta}_{3}$ 5 5.5 3.6 [Kushwaha2024] Alloys $\text{TaIr}_{2}\text{B}_{2}$ 3.6 2.7 5.2 [Gornicka2020] Borides $\text{Pt}_{1.5}\text{Ir}_{1.5}\text{Zr}_{5}$ 4.9 4.9 3.2 [Watanabe2023] Alloys $\text{Pt}_{1.2}\text{Ir}_{1.8}\text{Zr}_{5}$ 4.5 5.1 2.8 [Watanabe2023] Alloys $\text{Cu}_{2}\text{Si}$ 2.1 2.7 4.1 [Yan2019] Silicides HfRhGe 3.8 4 1.7 [P2024] Germanides $\text{Pt}_{2.7}\text{Ir}_{0.3}\text{Zr}_{5}$ 6 2.5 3.8 [Watanabe2023] Alloys $\text{Mo}_{20}\text{Re}_{6}\text{Si}_{4}$ 7.7 12.3 5.4 [Kaplan2025] Silicides $\text{Pt}_{2.1}\text{Ir}_{0.9}\text{Zr}_{5}$ 5.8 3.5 3.4 [Watanabe2023] Alloys $\text{Pt}_{0.9}\text{Ir}_{2.1}\text{Zr}_{5}$ 5.1 5.5 2.7 [Watanabe2023] Alloys $\text{Pt}_{2.4}\text{Ir}_{0.6}\text{Zr}_{5}$ 6.2 3.9 3.6 [Watanabe2023] Alloys HfIrGe 3.1 3.5 5.64 [Meena2025] Germanides $\text{Pt}_{0.45}\text{Ir}_{2.55}\text{Zr}_{5}$ 5.3 5.9 2.6 [Watanabe2023] Alloys $\text{Pt}_{0.6}\text{Ir}_{2.4}\text{Zr}_{5}$ 5.3 5.9 2.6 [Watanabe2023] Alloys $\text{NbIr}_{2}\text{B}_{2}$ 4.2 3.2 7.2 [Gornicka2020] Borides $\text{Pt}_{3}\text{Zr}_{5}$ 3.5 2.8 7.3 [Hamamoto2018] Alloys $\text{Re}_{8}\text{NbTa}$ 3.4 3.9 7.7[Kushwaha2024-1] Alloys ReTc 6.9 6.3 11.5 [Cerqueira2023] [^1] DFT predicted $T_{c}$ Alloys
V.2 Using the compounds in the Materials Project